OpenStudy (anonymous):

How do I prove this identity: cos^2x - cos^4x = cos^2xsin^2x? o.o

OpenStudy (lgbasallote):

well for one you can factor out cos^2x and see that you'll arrive with an identity

OpenStudy (anonymous):

maybe start with \[\cos^2(x)-\cos^4(x)=\cos^2(x)(1-\cos^2(x))\]

OpenStudy (anonymous):

what do you mean?

OpenStudy (lgbasallote):

let \(\cos^2 x = a\) \[\cos^2 x - \cos^4 x = a - a^2\] agree?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so let's factor out a \[a(1 - a)\] right?

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

let's substitute a back to \(\cos^2 x\) \[\cos^2x(1 - \cos^2 x)\]

OpenStudy (anonymous):

ohhh and then cos2x(1-sin2x)?

OpenStudy (anonymous):

oh no. boo

OpenStudy (lgbasallote):

uhmm no.... \(1 - \cos^2 x \ne 1 - \sin^2x\) however....let the identity \(\sin^2 x + \cos^2 x = 1\) be your guide

OpenStudy (anonymous):

yup i was looking at that

OpenStudy (anonymous):

so then what happens after cos2x(1−cos2x)

OpenStudy (lgbasallote):

try to make \("\sin^2 x + \cos^2 x = 1"\) look like \("1 - \cos^2 x"\) can you d that??

OpenStudy (anonymous):

isn't that for sin2x?

OpenStudy (lgbasallote):

yup!

OpenStudy (anonymous):

what do I do with it? o.o

OpenStudy (lgbasallote):

\[\cos^2 x (1 - \cos^2x) = \cos^2 x (\sin^2 x)\]

OpenStudy (lgbasallote):

you have just proven it :D

OpenStudy (anonymous):

ohhhhhhh. omg, i need to look at the right side too >< thank you!

OpenStudy (anonymous):

i keep forgetting

OpenStudy (lgbasallote):

hahaha but remember that in proving you can only touch ONE side..

OpenStudy (lgbasallote):

but your goal is to make it look like the other side